OpenStudy (anonymous):

rewrite with rational exponents square root (x^5)

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt{x^5}=\left(x^5\right)^{\frac{1}{2}}\] \[\LARGE \sqrt{x^5}=x^{5\times\frac{1}{2}}\] \[\LARGE\sqrt{x^5}=x^{\frac{5}{2}}\]

OpenStudy (anonymous):

how is it 1/2?

jimthompson5910 (jim_thompson5910):

taking the square root of any number is the same as raising that number to the one-half power

OpenStudy (anonymous):

oh okay